Brad Young is married with one withholding, earns $3,600 per month. Brad's employer pays him on a semimonthly basis. He has two pre-tax deductions: qualified health insurance deductions of $100 per pay and a 5% 401(k) contribution. Brad will pay $157 in federal taxes.

Required: What is Brad's net pay? Round calculations to 2 decimal places and share calculations.

Employee Name: Brad Young

Married

# of withholding: 1

Wages: ???

Gross: ???

Insurance: $100

401(k): 5%

Federal Taxable wages: ???

FICA taxable wages: ???

Federal Withholding: ???

Social Security withholding: ???

Medicare withholding: ???

Net Pay: ???
Answer #1

Employee name

Married/single

# of withhold

Wages

Gross

Insurance

401(k)

Federal taxable ages

FICA taxable ages

Federal WH

Social security WH

Medicare WH

Neet pay

Brad Young

married

1

$43200

$1800

100

90

1710

1800

157

111.60

26.10

$1315.30

Wages = 3600*12 = 43200

Gross = 3600/2 = 1800 (Semimonthly basis)

401(k) contribution = 1800*5% = 90

Federal taxable ages = 1800-90 = $1710

FICA tax – social security tax = 1800*6.2% = $356.50

FICA tax – Medicare tax = 1800*1.45% = $83.38

Net Pay = gross pay – social security tax – medicare tax – federal income tax withheld – health insurance deduction – contribution to retirement plan = 1800-111.60-26.10-157-100-90 = 1315.30